# A Particle-in-Cell Simulation Framework for Thomson Scattering Analysis in Inertial Confinement Fusion

### v1

@@ 4.3 Thomson scattering from perturbations with a wave vector deviated from matching condition
where $L_{\rm pert}=700\,c/\omega_{0}$ is the length of the perturbation region, and $x_{0}=980\,c/\omega_{0}$ represents its center in the simulation domain. We set the current to zero beyond the width $L_{pert}$. Figures 8 show the spatial envelope of $\vec{J}{\rm pert}$ along with the resulting electromagnetic field for two cases: matched, with $\omega_{\rm pert}/k_{\rm pert}=c$, and mismatched, with $\omega_{\rm pert}/k_{\rm pert}=2c$.
In the matched case, a propagating electromagnetic wave is excited within the perturbed region and radiates outward. By contrast, the mismatched case features a near-field electromagnetic perturbation localized to the driven region; at the boundary, scattered light at frequency $\omega_{pert}$ is emitted into the surrounding vacuum. These results indicate that scattered light can still be generated even under unmatched wave-vector.
Since driven perturbations that deviate from the matching condition can still produce SCTS signals, the actual ion modes underlying the diagnosed SCTS signals should not be restricted to the exact matching condition. This needs to be carefully investigated and taken care of in analyzing our experimental results on Shenguang facility.
@@ 5 Summary
This work was supported by the National Key R&D Program of China, No. 2023YFA1608400 and National Natural Science Foundation of China (NSFC) (Grant No. 12275269, U2430207, 12275251). The numerical simulations in this paper were conducted on Hefei advanced computing center.

### v2

@@ 4.3 Thomson scattering from perturbations with a wave vector deviated from matching condition
where $L_{\rm pert}=700\,c/\omega_{0}$ is the length of the perturbation region, and $x_{0}=980\,c/\omega_{0}$ represents its center in the simulation domain. We set the current to zero beyond the width $L_{pert}$.
This 1D model is a simplified setup to explicitly reveal the emission mechanism. The scenarios map directly with the 2D simulations: the imposed current $J_{\mathrm{pert}}$ in 1D represents the source from the probe–perturbation interaction in 2D. The mismatch in Fig. 8(a-d) corresponds to choosing $\omega_{\mathrm{pert}}/k_{\mathrm{pert}}\neq c$ (i.e., violating the EM dispersion relation), analogous to the $\Delta\vec{k}$ mismatch in Fig. 7(b). Please note that the amplitude of $\vec{J}_{\rm pert}$ is not crucial to the primary aim of the simulations presented in Fig. (8), which is to demonstrate whether emission occurs under matched or mismatched conditions.
Figure (8) shows the spatial envelope of $\vec{J}_{\rm pert}$ and the resulting electromagnetic field for five cases with $\omega_{\rm pert}=\omega_{0}$ and $k_{\rm pert}$ = 0.5, 0.8, 1.5, 1.2, 1.0 $\omega_{0}/c$. In the case corresponding to matched wave vector ($k_{\rm pert}=1.0$), a propagating electromagnetic wave is excited within the perturbed region and radiates outward [Fig. 8(e)]. By contrast, other cases with mismatched wave vectors feature a near-field electromagnetic perturbation localized to the driven region; at the boundary, scattered light at frequency $\omega_{pert}$ is emitted into the surrounding vacuum. These results indicate that scattered light can still be generated even under unmatched wave-vector. Moreover, We observe that the intensity of the scattered light is highest when $\omega_{\mathrm{pert}}/k_{\mathrm{pert}}$ approaches the light speed $c$, corresponding to the matched condition. The emission amplitude decreases systematically as the ratio deviates from $c$, consistent with the expected behavior for off-resonant driving.
Additionally, we point out that the primary purpose of the vacuum simulation is to isolate and illustrate the core physical mechanism of eigenmode excitation by a localized non-eigenmode source, not to model the full plasma dispersion quantitatively. While this simplification will inevitably lead to quantitative deviations in the wave properties (e.g., a slight shift in the effective wave number for a given frequency), it does not alter the fundamental process by which a non-propagating field excites a propagating eigenmode at a boundary.
Therefore, since driven perturbations that deviate from the matching condition can still produce SCTS signals, the actual ion modes underlying the diagnosed SCTS signals should not be restricted to the exact matching condition. This needs to be carefully investigated and taken care of in analyzing our experimental results on Shenguang facility.
@@ 5 Summary
This work was supported by the National Key R&D Program of China, No. 2023YFA1608400 and National Natural Science Foundation of China (NSFC) (Grant No. 12275269, U2430207, 12275251). The numerical simulations in this paper were conducted on Hefei advanced computing center. We thank the UCLA-IST OSIRIS Consortium for the use of OSIRIS.
Appendix A Influence of Particle-per-Cell Resolution on Signal-to-Noise Ratio
To evaluate the effect of particle-per-cell (PPC) resolution on the thermal collective Thomson scattering spectra, we performed PIC simulations with 25, 100, and 400 macroparticles per cell for both electrons and ions. All other parameters were identical to those used in Fig. 3(a). The resulting spectra at a scattering angle $\theta_{s}=90^{\circ}$ are shown in Fig. 9(a-c).
Figure 9: Simulated (blue dashed line) and theoretical (red solid line) scattering spectra at $\theta_{s}=90^{\circ}$ along with the correlation coefficient $\rho$ between them. (a–c): single simulations with domain size $200c/\omega_{0}\times 200c/\omega_{0}$ and PPC = 25, 100, and 400, respectively. (d): single simulation with enlarged domain $800c/\omega_{0}\times 800c/\omega_{0}$ and PPC = 100. (e): ensemble-averaged spectra over 16 independent simulations with domain $200c/\omega_{0}\times 200c/\omega_{0}$ and PPC = 100.
The absolute amplitude of the simulated CTS signal depends on PPC because the fluctuation level in PIC simulations is influenced by finite macroparticle statistics. Increasing PPC reduces the discrete particle noise, leading to a decrease in both the background fluctuation level and the scattered spectral amplitude. Therefore, the absolute signal intensity from an individual simulation is not a strictly converged quantity.
In contrast, the normalized spectral shape—which is the primary diagnostic observable—remains consistent across the tested PPC values. The characteristic double-peak structure is preserved, and the agreement with the theoretical spectrum, quantified by the correlation coefficient $\rho$ (Fig. 9(a–c)), changes only marginally with increasing PPC. Although higher PPC reduces statistical noise, a single simulation remains a stochastic realization with finite sampling of fluctuation modes, and thus the ensemble-averaged spectrum is not fully represented.
Improved convergence is achieved either by enlarging the simulation domain or by ensemble averaging over independent runs. As shown in Fig. 9(d) and (e), both a domain enlarged by a factor of 16 and the average of 16 independent simulations yield correlation coefficients exceeding 0.99.
We therefore employ ensemble averaging to obtain high-fidelity CTS spectra within the adopted simulation framework, while a quantitative comparison of absolute spectral intensities with experimental measurements is beyond the scope of this work.
Appendix B Influence of Wave Vector Broadening on Angular Broadening of Scattered Light
The angular distribution of the scattered light shown in Fig. 6(b) exhibits finite peak widths for both the probe beam and the scattered signal. To quantify the role of wave-vector broadening in determining the angular spread of the scattered light, we analyze the dependence of the scattered angular width on the spectral broadening of both the probe beam and the driven density perturbation. As shown in Fig. 10(a), the half-width at half-maximum (HWHM) of the scattered light exceeds that of the probe beam, indicating that the wave-vector broadening of the density perturbation must also be taken into account.
Figure 10: (a) Half-width at half-maximum (HWHM) of the angular distribution as a function of scattering angle (blue asterisks). The gray dashed line denotes the HWHM of the probe beam. Red circles denote the quadrature sum of the probe-beam HWHM and the HWHM associated with wave-vector broadening of the density perturbations. (b) HWHM of wave vector spectrum of electron density perturbation corresponding to different scattered light.
In Fig. 10(b), the wave-vector broadening $\Delta k_{\rm pert}$ of the density perturbation at different $k_{y}$ is evaluated using the HWHM of the spectral peak. After considering the widths of both the probe beam and density modulations via $\sqrt{\Delta\theta_{probe}^{2}+\Delta\theta_{pert}^{2}}$ with $\Delta\theta_{pert}=\frac{\Delta k_{pert}}{k_{i}}$, the total contributions are plotted in Fig. 10(a) by red circles. We find that these results cannot explain the growing width of the scattered lights for greater $\theta_{s}$. These results demonstrate that although wave-vector broadening contributes to the angular spread of the scattered signal, it is insufficient to fully account for the scattering observed under wave-vector-mismatched conditions.
Appendix C Influence of Wave Vector Broadening on Mismatched Scattering
In realistic Thomson scattering configurations, the interaction occurs within a finite spatial region of characteristic length $L$, leading to a finite wave-vector broadening of order $1/L$. Under wave-vector-mismatched conditions, such finite spectral widths of the probe beam and the driven density perturbation could in principle allow the scattering to happen if the imposed wave vector deviation lies within the effective spectral width.
Figure 11: (a) Half width at half maximum (HWHM) of the quadrature sum of the probe-beam and density-perturbation broadenings along with the wave vector deviation from matched condition $\Delta k$. (b–c) Scattered spectra at $40.5^{\circ}$ for scaling factors of 0.75 and 1.5, respectively.
To test whether the observed mismatched scattering can be attributed solely to finite spectral broadening, we varied the waist of both the probe and drive beams in the simulation corresponding to Fig. 7, using scaling factors of 0.75 and 1.5 relative to the reference case. The corresponding wave-vector broadenings and scattered spectra are shown in Fig. 11.
To evaluate the combined effect of these broadenings on the scattering process, we estimate an effective spectral width given by the quadrature sum $\sqrt{{\Delta k_{probe}}^{2}+{\Delta k_{pert}}^{2}}$. As shown in Fig. 11(a), even in the smallest-waist case, the combined HWHM (blue asterisks) and HWTM (red circles) remain substantially smaller than $|\Delta k|$ (gray dashed line). Nevertheless, clear scattering persists in all cases, as shown in Fig. 11(b) and (c), indicating that this scattering under wave vector mismatched condition cannot be fully explained by the finite wave vector broadening.